Perform the indicated opeations.
Simplify:
-9 to the -2power

Answer by stanbon(48569)

(-9)^(-2) = 81^(-1) = 1/81
OR
-9^(-2) = -[81^-1] = -1/81
Comment: The answer depends on whether you
mean - 9^-2 or wheter you mean (-9)^-2

==============================
Cheers,
Stan H.